OpenStudy (anonymous):

Which of the following shows a proportion? four over nine equals eighteen over eight three over eight equals six over four six over twelve equals eight over four nine over fifteen equals twelve over twenty.

OpenStudy (anonymous):

van somebody please help me

OpenStudy (rsadhvika):

cross multiply and see which option is true

OpenStudy (rsadhvika):

\(\large \frac{4}{9} = \frac{18}{8}\) cross multiply \(\large 4\times 8 = 9 \times 18\) \(\large 32 = 162\) false

OpenStudy (rsadhvika):

similarly try other options and see for which you get true

OpenStudy (jkristia):

yeah, cross multiply or simply check if either side is the same ration, and if not, the statement is not true. clearly \(\frac{4}{9} \neq \frac{18}{8} \) so this statement is not true

OpenStudy (texaschic101):

3/8 = 6/4 (3)(4) = (6)(8) 12 = 48 <-- incorrect 6/12 = 8/4 (6)(4) = (8)(12) 24 = 96 < -- incorrect 9/15 = 12/20 (9)(20) = (12)(15) 180 = 180 < -- correct .....this is your answer :)